Assume that cigarettes cost $7 per pack and consider a 21 year old college student smoker who smokes 15 packs of cigarettes per month. If the student quits smoking and each month invests the amount she would have spent on cigarettes in a savings plan that averages a 4% annual return, how much will she have saved by the time she is 65?

Answer by Theo(13342)

cigarettes cost 7 collars a pack.
smoker smokes 15 packs a month.
15 * 7 = 105 dollars per month spent on cigarettes.

invest that 105 dollars at the end of each month for 65 - 21 = 44 years at 4% annual return will yield a future value as shown below.

use the online calculator at https://arachnoid.com/finance/

here's a display of the use of that calculator to solve this problem.

$$$

your inputs to this calculator would be:

present value = 0
future value = 0
number of time periods = 44 years * 12 = 528 months.
payment per month = 15 * 7 = 105
interest rate per month = 4% per year / 12 = .333333.....%.
payment is made at the end of each month.

you would then click on FV and the calculator tells you that you would have 155,056.68 dollars when you're 65 years old.

528 * 105 = 55,440 of that would be money you invested over the 44 year period.

the rest is interest earned from those payments.